OpenStudy (anonymous):

Consider the reaction Mg(s) + I2(s) -> MgI2(s) Identify the limiting reagent if there are 100 atoms of Mg and 100 molecules of I2

OpenStudy (anonymous):

We need to find which reactant, Mg or I, will make the smallest number of moles of MgI2.

OpenStudy (anonymous):

Let's start with Mg.

OpenStudy (anonymous):

First, turn the 100 atoms of Mg into moles of Mg. Let me know if you need help to do that.

OpenStudy (anonymous):

That's the 6.022*10^23 thing right?

OpenStudy (anonymous):

Yes. \[100 atoms Mg \times \frac{ 1 mol }{ 6.02^{23} atoms Mg} \]

OpenStudy (anonymous):

Okay done with that, got 1.66*10^-22 mol

OpenStudy (anonymous):

My bad for taking so long. I was checking

OpenStudy (anonymous):

It's fine. Go on:)

OpenStudy (anonymous):

Now, see how many moles of MgI2 you get. Let me know if you need help with that.

OpenStudy (anonymous):

Do the same for I2

OpenStudy (anonymous):

Yeah, I'm having trouble with that.

OpenStudy (anonymous):

Alright, the equation states that there is a 1:1 relationship with Mg and MgI2

OpenStudy (anonymous):

It will look like this, \[1.66^{-22} moles of Mg \times \frac{ 1 mole MgI2 }{ 1 mole Mg} = \]

OpenStudy (anonymous):

So it will be \[1.66^{-22} moles of MgI2\]

OpenStudy (anonymous):

Do the same for I2. Let me know if you need help with that.

OpenStudy (anonymous):

What's different in finding the I2 in mol since it's molecules not atoms?

OpenStudy (anonymous):

Not much. Remember\[1 mole = 6.02^{23} atoms, molecules, formula units, etc\]

OpenStudy (anonymous):

Oh okay. I'll find that then.

OpenStudy (anonymous):

Wait shouldn't it just be the same thing then?

OpenStudy (anonymous):

So try it like this \[100 molecules of I _{2} \times \frac{ 1 mole of I2 }{ 6.02^{23} molecules of I2} = \]

OpenStudy (anonymous):

Yes, but you have to do the boring stuff like the rest of us

OpenStudy (anonymous):

Did that. Agreed, it is boring.

OpenStudy (anonymous):

It's the same thing as the Mg

OpenStudy (anonymous):

Yep. So what now?

OpenStudy (anonymous):

Was there any more information?

OpenStudy (anonymous):

Ummm... it just said to find the limiting regent for it. That's all.

OpenStudy (anonymous):

Ok, I say we figure which reactant has the smallest mass then.

OpenStudy (anonymous):

Okay then. That's the Mg then.

OpenStudy (anonymous):

I would go with that.

OpenStudy (anonymous):

I hope I didn't waste your time. Good luck.

OpenStudy (anonymous):

No but what do I do with that information? What's the regent? Sorry, I'm pretty new to this. It's the summer assignment.

OpenStudy (anonymous):

That's cool. I say which ever has the smallest mass will run out first and thus, is the limiting reactant.

OpenStudy (anonymous):

I could be wrong

OpenStudy (anonymous):

It's all I have right now. I can check the answer but I still don't know how to get it. Thanks, still.

OpenStudy (anonymous):

It just says "stoichiometric mixture"

OpenStudy (anonymous):

@asib1214 what do you think?

OpenStudy (anonymous):

I googled that, and it said that a stoichiometric mixture is a reaction with no excess reactants. In other words, there is no limiting reactant. Sorry about the drawn out process.

OpenStudy (anonymous):

It's fine. I at least somewhat understand it now. It's a start. Thanks so much!

OpenStudy (anonymous):

No prob. Thank you too

OpenStudy (rogue):

There actually isn't a limiting reagent in this case unless the question is worded incorrectly.

OpenStudy (rogue):

According to the reaction Mg (s) + I2 (s) -> MgI2 (s), one magnesium atom reacts with every one I2 molecule to produce 1 molecule of MgI2. Since Mg and I2 react in a 1:1 ratio, you simply need equal amounts to Mg and I2 to not have any shortage of reactants. In this question, you do have equal amounts: 100 atoms of Mg for 100 molecules of I2. There is no excess or shortage—all of the Mg and I2 get used up.

OpenStudy (vincent-lyon.fr):

@Rogue is right Simply look at the stoichiometric numbers : you can read them as moles or as microscopic units (atoms, molecules, ions...). Since the numbers are 1 and 1, then 100 and 100 represent a stoichiometric mixture.